# Aligning Human and LLM Judgments: Insights from EvalAssist on Task-Specific Evaluations and AI-assisted Assessment Strategy Preferences

### v1

@@ Abstract.
Evaluation of large language model (LLM) outputs requires users to make critical judgments about the best outputs across various configurations. This process is costly and takes time given the large amounts of data. LLMs are increasingly used as evaluators to filter training data, evaluate model performance or assist human evaluators with detailed assessments. To support this process, effective front-end tools are critical for evaluation. Two common approaches for using LLMs as evaluators are direct assessment and pairwise comparison. In our study with machine learning practitioners (n=15), each completing 6 tasks yielding 131 evaluations, we explore how task-related factors and assessment strategies influence criteria refinement and user perceptions. Findings show that users performed more evaluations with direct assessment by making criteria task-specific, modifying judgments, and changing the evaluator model. We conclude with recommendations for how systems can better support interactions in LLM-assisted evaluations.
@@
journalyear: 2024
@@ 1. Introduction
Large language models are foundation models that can be used for a variety of tasks such as summarization, text generation, concept extraction, analysis, or classification. Benchmarks such as Helm(Liang et al., 2022), BigBench(Ghazal et al., 2013), or MMLU(Hendrycks et al., 2020) can provide guidance in what language model to pick for a certain task. However, in practice, they are insufficient when it comes to specific use cases, use case specific data, or creative tasks (Zheng et al., 2024a). Often human evaluation is used to assess model fitness for specific use cases and data but it is costly and takes time given the large volumes of data being generated. Practitioners are increasingly using large language models also as evaluators of the output of large language models across various contexts including filtering training data, evaluating model performance, assessing prompt effectiveness, and assisting human evaluators with detailed assessments and explanations (Huang et al., 2024; Chen et al., 2024; Pan et al., 2024; Raju et al., 2024). This approach is often referred to as LLM-as-a-judge. [^1] Note that, in this paper, we use the terms judge and evaluator interchangeably.
@@
However, relying solely on LLM evaluators is not without risks; as with many tasks, evaluator models can also hallucinate or provide explanations that lack coherence, underscoring the necessity of keeping humans in the loop. Although LLMs may not always be accurate, they can potentially reduce workload by flagging outputs that require human input due to low confidence. Additionally, LLM evaluators also offer a lot of flexibility because practitioners can customize evaluation criteria — such as conciseness, faithfulness to a context, conversational naturalness, or succinctness — enabling more targeted and effective assessments tailored to specific use cases. Many tools have emerged to help users create criteria to evaluate their outputs. The process of criteria creation for evaluation is often iterative, and the concept of ”criteria drift” may arise. While predefined criteria may help users to assess outputs, in practice, the act of grading also helps users refine and redefine these criteria.
@@
Two predominant forms of AI-assisted assessments with LLMs have emerged: Direct Assessment and Pairwise Assessment. Direct Assessment scores specific criteria (often part of a rubric) to evaluate whether outputs meet the criteria, while Pairwise Assessment compares pairs of outputs against broader, high-level criteria. Each method has its own strengths and limitations, which our study explores by examining how they affect user interaction with both the criteria and the evaluation process. To support this investigation, we developed EvalAssist, a system that allows users to view outputs and iteratively refine their evaluation criteria. Our focus was on understanding how users adjust their criteria, the changes they make, and how they ultimately find satisfaction with both the criteria and the evaluation process. To understand the how users develop and refine criteria to achieve alignment with LLM evaluators, we ran a within-subject study with 15 practitioners (data scientists, software engineers, and AI engineers) who have been involved in model performance projects. In our study we pose the following research questions:
  • RQ1: What do practitioners prioritize in evaluation criteria when using LLMs as judges, and how do their priorities differ?
  • RQ2: What strategies do users employ to refine their criteria for achieving human-AI alignment in LLM-as-a-judge evaluations?
@@
  • RQ3: How do task-related factors and judge strategy impact how practitioners refine criteria?
@@
    • RQ3A How do task and judge strategy influence the total number of evaluations performed?
    • RQ3B How do task and form of assessment (direct vs. pairwise) affect the degree of human-AI alignment?
@@
  • RQ4: How do task-related factors and judge strategy impact user perceptions of the judge?
    • How do they affect users’ trust in AI?
    • How do they shape users’ perception of positional bias?
    • How do they influence users’ perception of explanations provided by the judge?
@@
    • How do they impact users’ cognitive load during the evaluation process?
  • RQ5: Which assessment strategy do users prefer?
@@
  • We introduce EvalAssist, a tool designed to help practitioners refine evaluation criteria through both direct and pairwise assessment strategies. It provides positional bias metrics from the AI judge, as well as an explanation for each judgment.
  • We present results from within-subjects controlled experiment with machine learning practitioners (n=15) providing insights into how they refine evaluation criteria and uncover key differences between the two evaluation strategies.
  • Based on our findings, we offer design suggestions for AI-assisted evaluation systems.
Our findings show that users conduct more evaluations under the direct assessment condition. Users refine criteria in multiple ways, such as making criteria more specific or general, adjusting their own judgments, or modifying the AI evaluator’s outputs. Explanations are perceived as more helpful in the direct assessment condition. Users prefer Direct Assessment when they need clarity and control over individual item evaluations, and Pairwise Assessment when evaluating nuanced or subjective criteria.
@@ 2.1. Human AI Collaboration in AI-assisted Evaluation
Recently, several AI-assisted evaluation tools have emerged, with varying focuses including: improving the iterative nature of prompt refinement or refining criteria for evaluation. Across these systems, the evolving nature of user-defined criteria is emphasized, acknowledging that such criteria are not static but adapt in response to AI outputs and user feedback. This iterative process is fundamental in tools that support criteria refinement and prompt adjustments, reinforcing the dynamic interaction between humans and AI in the evaluation process. A prominent aspect shared among these tools is the role of human-in-the-loop evaluation. Systems like EvalGen (Shankar et al., 2024), ChainForge (Arawjo et al., 2024), EvalLM (Kim et al., 2023a), and LLM Comparator (Kahng et al., 2024) integrate human feedback as a key element of the evaluation loop. While AI systems assist in generating evaluations, they rely on human judgment to ensure alignment with user preferences. One of the key challenges in this interactive process is criteria drift (Shankar et al., 2024), where users adjust their evaluation standards as they encounter new outputs. Prior systems illustrate this behavior, as users often modify their criteria after receiving AI-generated responses that deviate from initial expectations. This flexibility is critical, highlighting the need for evaluation systems that allow criteria adjustments throughout the evaluation process, rather than imposing rigid, predefined standards.
@@ 2.1.1. Criteria Iteration
Crafting effective criteria typically requires multiple iterations. Criteria components such as name, definition, scale, and examples often need definition and refinement as users evaluate outputs. Related work (Kim et al., 2023b) indicates that users often develop new criteria during evaluations. To facilitate this, EvalAssist includes a real-time feedback system that allows users to immediately see the impact of criteria modifications. Unlike systems that rely heavily on predefined metrics or expert-labeled data, EvalAssist enables users to define evaluation criteria in natural language and to iteratively refine these criteria based on feedback from the AI model. Unlike other AI-assisted evaluation tools that combine prompting engineering with criteria definition, EvalAssist simplifies the LLM-as-a-judge process by allowing users to focus solely on defining criteria. This approach recognizes that developers often rely on external workflows to adjust configurations (e.g., model temperature) and experiment with different models and prompts to generate responses (Desmond et al., 2024).
@@ 2.1.2. Visualization
Other tools focus on providing intuitive, interactive interfaces that facilitate complex evaluation tasks. The use of interactive and visual interfaces is another notable feature across these tools. Allowing users to visually compare model outputs in real-time, provides comprehensible evaluation experience (Kahng et al., 2024). Many existing tools share common themes such as iterative refinement, user-centered evaluation, and scalability. Together, they reflect a growing trend in human-AI collaboration, aiming to create more flexible, subjective, and adaptable evaluation systems that effectively combine human insight with AI capabilities.
@@ 2.1.3. Addressing Bias in AI-Assisted Evaluation
LLM Evaluators, like their human counterparts, exhibit biases. These biases include but are not limited to: positional bias, which is when judges consistently favor one side of a pair, regardless of the actual quality of the answers, self-enhancement bias when a model prefers its own responses, and verbosity bias when an LLM judge favors longer responses even if they are not a better alternative (Zheng et al., 2024a). Many of the existing AI-assisted tools do not flag these kinds of biases to users. Considering the persistent challenge of bias, systems should both provide transparency when bias occurs and implement bias mitigation strategies that include swapping answer order to reduce position bias (Zheng et al., 2024a) and treating inconsistent results as ties, or by randomly assigning positions in large datasets (Li et al., 2023) (Zheng et al., 2024a). EvalAssist includes a check for positional bias and indicates whether positional bias exists.
@@
Two of the most common judgment strategies in evaluation are direct assessment and pairwise comparison. Direct assessment involves outputting a scalar indicator of quality (e.g., assigning a score or rating to an item) (Zheng et al., 2024a), while pairwise comparison determines which of two outputs is preferred based on specific criteria. Both approaches have advantages and disadvantages depending on the context and task. One limitation of pairwise comparison is scalability. As the number of items to be evaluated increases, the number of required comparisons grows quadratically, making this method less feasible for large-scale evaluations. However, pairwise comparisons can be more effective at identifying subtle differences between outputs and according to prior research is an easier task for both humans and LLMs compared to rating a single output, often yielding higher accuracy in LLM-as-a-judge benchmarks (Bai et al., 2024; Gehrmann et al., 2023; Li et al., 2019; Zheng et al., 2024b). In contrast, direct assessment can efficiently evaluate multiple items at once, but it may struggle to detect fine distinctions between outputs.
Direct assessment often utilizes rubrics with multiple dimensions, while pairwise comparison may focus on a single dimension of preference. Additionally, the nature of the criteria—whether objective (e.g., grammatical accuracy) or subjective (e.g., creativity or tone)—plays a significant role in shaping the evaluation process. Prior work has explored LLMs’ ability to make selections based on user-defined preferences across a wide range of criteria (Chiang and Lee, 2023a, b; Zheng et al., 2024a). For example, prior work investigated how LLMs handle subjective judgments across dimensions such as brevity, formality, honesty, creativity, and political tone, highlighting the variability that emerges when criteria are subjective. EvalAssist allows users to select the evaluation strategy that best fits the task.
@@
3. EvalAssist

### v2

# EvalAssist: Insights on Task-Specific Evaluations and AI-Assisted Judgment Strategy Preferences

@@ Abstract.
With the broad availability of large language models and their ability to generate vast outputs using varied prompts and configurations, determining the best output for a given task requires an intensive evaluation process, one where machine learning practitioners must decide how to assess the outputs and then carefully carry out the evaluation. This process is both time-consuming and costly. As practitioners work with an increasing number of models, they must now evaluate outputs to determine which model performs best for a given task. LLMs are increasingly used as evaluators to filter training data, evaluate model performance or assist human evaluators with detailed assessments. Our application, EvalAssist, supports this process by aiding users in interactively refining evaluation criteria. In our study with machine learning practitioners (n=15), each completing 6 tasks yielding 131 evaluations, we explore how task-related factors and judgment strategies influence criteria refinement and user perceptions. Findings show that users performed more evaluations with direct assessment by making criteria task-specific, modifying judgments, and changing the AI evaluator model. We conclude with recommendations for how systems can better support practitioners with AI-assisted evaluations.
@@
journalyear: 2025copyright: ccconference: The 38th Annual ACM Symposium on User Interface Software and Technology; September 28-October 1, 2025; Busan, Republic of Koreabooktitle: The 38th Annual ACM Symposium on User Interface Software and Technology (UIST ’25), September 28-October 1, 2025, Busan, Republic of Koreadoi: 10.1145/3746059.3747740isbn: 979-8-4007-2037-6/2025/09
@@ 1. Introduction
Large language models (LLMs) are foundation models that can be used for a variety of tasks such as summarization, text generation, concept extraction, analysis, or classification. Benchmarks such as Helm(Liang et al., 2022), BigBench(Ghazal et al., 2013), or MMLU(Hendrycks et al., 2020) can provide guidance in what language model to pick for a certain task. However, in practice, they are insufficient when it comes to specific use cases, use case specific data, or creative tasks (Zheng et al., 2024). In enterprise contexts, where LLMs are increasingly deployed (Desmond and Brachman, 2024; Brachman et al., 2024), machine learning practitioners have relied on ad hoc, manual evaluation methods to assess model outputs against diverse and evolving criteria (Desmond et al., 2024). Such approaches are resource-intensive and challenging to scale. Recently, the use of LLMs as evaluators of outputs, referred to as “LLM-as-a-judge” (Desmond et al., 2024; Pan et al., 2024) gained traction. This approach shows promise across contexts including filtering training data, evaluating model performance, assessing prompt effectiveness, and assisting human evaluators with detailed assessments and explanations (Huang et al., 2024; Chen et al., 2024; Pan et al., 2024; Raju et al., 2024). [^1] Note that, in this paper, we use the terms judge and evaluator interchangeably.
@@
However, relying solely on AI Evaluators is not without risks; as with many tasks, evaluator models can also hallucinate or provide explanations that lack coherence, underscoring the necessity of keeping humans in the loop. Although LLMs may not always be accurate, they can potentially reduce workload by flagging outputs that require human input due to low confidence. Additionally, AI Evaluators also offer a lot of flexibility because practitioners can customize evaluation criteria, such as conciseness, faithfulness to a context, conversational naturalness, or succinctness, enabling more targeted and effective assessments tailored to specific use cases. Many tools have emerged to help users create criteria to evaluate their outputs. The process of criteria creation for evaluation is often iterative, and the concept of ”criteria drift” may arise (Shankar et al., 2024). While predefined criteria may help users to assess outputs, in practice, the act of grading also helps users refine and redefine these criteria.
@@
Two predominant forms of AI-assisted assessments with LLMs have emerged: direct assessment and pairwise comparison. We will refer to these as judgment strategies throughout the paper. Direct assessment scores specific criteria (often part of a rubric) to evaluate whether outputs meet the criteria, while pairwise comparison compares pairs of outputs against broader, high-level criteria. Each method has its own strengths and limitations, which our study explores by examining how they affect user interaction with both the criteria and the evaluation process. To support this investigation, we developed EvalAssist, a system that allows users to view outputs and iteratively refine their evaluation criteria. Our focus was on understanding how users adjust their criteria, the changes they make, and how they ultimately find satisfaction with both the criteria and the evaluation process. To understand how users develop and refine criteria to achieve alignment with AI Evaluators, we ran a within-subject study with 15 practitioners (data scientists, software engineers, and AI engineers) who have been involved in model performance projects. In our study we pose the following research questions:
  • RQ1: How do machine learning practitioners approach the development and refinement of evaluation criteria when using LLM-as-a-judge, and what strategies, characteristics, and evaluative dimensions emerge in their criteria creation process?
@@
  • RQ2: How do task-related factors and judge strategy impact how practitioners refine criteria?
@@
    • RQ2A How do task and judge strategy (direct vs. pairwise) influence the total number of evaluations performed?
    • RQ2B How do task and judge strategy affect the degree of human-AI alignment?
@@
  • RQ3: How do task-related factors and judge strategy impact user perceptions of the judge?
    • RQ3A How do they affect users’ trust in AI?
    • RQ3B How do they shape users’ perception of positional bias?
    • RQ3C How do they influence users’ perception of explanations provided by the judge?
@@
    • RQ3D How do they impact users’ cognitive load during the evaluation process?
  • RQ4: Which judgment strategy do users prefer?
@@
  • We introduce EvalAssist, an application designed to assist machine learning practitioners refine evaluation criteria through both direct and pairwise judgment strategies. It incorporates features such as positional bias metrics from the AI judge, explanations for each judgment, and the ability to select between two of the most common judgment strategies: pairwise comparison and direct assessment. By isolating generation from evaluation, enabling cost-efficient workflows, and supporting multiple AI evaluators, EvalAssist addresses key gaps in existing tools and offers a scalable solution for practitioners working with large datasets.
  • We present results from a within-subjects controlled experiment with machine learning practitioners (n=15) providing insights into how they refine evaluation criteria and uncover key differences between the two evaluation strategies.
  • Based on our findings, we offer design suggestions for AI-assisted evaluation systems. These include:
    • Exposing users to diverse task contexts to avoid over-specificity in criteria development.
    • Incorporating adaptive evaluation strategies that allow users to switch between direct assessment and pairwise comparison based on task complexity.
    • Enhancing explanation visibility, especially in pairwise conditions, by making explanations more concise and accessible.
    • Integrating a suite of bias indicators to increase transparency and user trust.
Our findings show that users conduct more evaluations under the direct assessment condition. Users refine criteria in multiple ways, such as making criteria more specific or general, adjusting their own judgments, or modifying the AI evaluator’s outputs. Explanations are perceived as more helpful in the direct assessment condition. Users prefer direct assessment when they need clarity and control over individual item evaluations, and pairwise comparison when evaluating nuanced or subjective criteria.
@@ 2.1. Human AI Collaboration in AI-assisted Evaluation
Recently, several AI-assisted evaluation tools have emerged, with varying focuses including: improving the iterative nature of prompt refinement or refining criteria for evaluation. Across these systems, the evolving nature of user-defined criteria is emphasized, acknowledging that such criteria are not static but adapt in response to AI outputs and user feedback. This iterative process is fundamental in tools that support criteria refinement and prompt adjustments, reinforcing the dynamic interaction between humans and AI in the evaluation process. A prominent aspect shared among these tools is the role of human-in-the-loop evaluation. Systems like EvalGen (Shankar et al., 2024), ChainForge (Arawjo et al., 2024), EvalLM (Kim et al., 2024a), and LLM Comparator (Kahng et al., 2024) integrate human feedback as a key element of the evaluation loop. While AI systems assist in generating evaluations, they rely on human judgment to ensure alignment with user preferences. One of the key challenges in this interactive process is criteria drift (Shankar et al., 2024), where users adjust their evaluation standards as they encounter new outputs. Prior research has shown that users often modify their criteria after receiving AI-generated responses that deviate from initial expectations (Kim et al., 2024a). This flexibility is critical, highlighting the need for evaluation systems that allow criteria adjustments throughout the evaluation process, rather than imposing rigid, predefined standards.
@@ 2.1.1. Criteria Iteration
Crafting effective criteria typically requires multiple iterations. Criteria components such as name, definition, scale, and examples often need definition and refinement as users evaluate outputs. Related work (Kim et al., 2024a) indicates that users often develop new criteria during evaluations. To facilitate this, EvalAssist includes a real-time feedback system that allows users to immediately see the impact of criteria modifications. Unlike systems that rely heavily on predefined metrics or expert-labeled data, EvalAssist enables users to define evaluation criteria in natural language and to iteratively refine these criteria based on feedback from the AI model. Unlike other AI-assisted evaluation tools that combine prompting engineering with criteria definition, EvalAssist simplifies the LLM-as-a-judge process by allowing users to focus solely on defining criteria. This approach recognizes that developers often rely on external workflows to adjust configurations like model temperature and experiment with different models and prompts to generate responses (Desmond et al., 2024).
@@ 2.1.2. Visualization
Other tools focus on providing intuitive, interactive interfaces that facilitate complex evaluation tasks. The use of interactive and visual interfaces is another notable feature across these tools. Allowing users to visually compare model outputs in real-time, provides comprehensible evaluation experience (Kahng et al., 2024). Many existing tools share common themes such as iterative refinement, user-centered evaluation, and scalability (Shankar et al., 2024; Arawjo et al., 2024; Kim et al., 2024a). Together, they reflect a growing trend in human-AI collaboration, aiming to create more flexible, subjective, and adaptable evaluation systems that effectively combine human insight with AI capabilities.
@@ 2.1.3. Addressing Bias in AI-Assisted Evaluation
AI Evaluators, like their human counterparts, exhibit biases. These biases include but are not limited to: positional bias, which is when judges consistently favor one side of a pair, regardless of the actual quality of the answers, self-enhancement bias when a model prefers its own responses, and verbosity bias when an LLM judge favors longer responses even if they are not a better alternative (Zheng et al., 2024). Many of the existing AI-assisted tools do not flag these kinds of biases to users. Considering the persistent challenge of bias, systems should both provide transparency when bias occurs and implement bias mitigation strategies that include swapping answer order to reduce position bias (Zheng et al., 2024) and treating inconsistent results as ties, or by randomly assigning positions in large datasets (Li et al., 2023) (Zheng et al., 2024). EvalAssist includes a check for positional bias and indicates whether positional bias exists.
@@
Two of the most common judgment strategies in evaluation are direct assessment and pairwise comparison (Zheng et al., 2024). Direct assessment involves outputting a scalar indicator of quality like assigning a score or rating to an item (Zheng et al., 2024), while pairwise comparison determines which of two outputs is preferred based on specific criteria. Both approaches have advantages and disadvantages depending on the context and task. One limitation of pairwise comparison is scalability. As the number of items to be evaluated increases, the number of required comparisons grows quadratically, making this method less feasible for large-scale evaluations. However, pairwise comparisons can be more effective at identifying subtle differences between outputs and, according to prior research, is an easier task for both humans and LLMs compared to rating a single output, often yielding higher accuracy in LLM-as-a-judge benchmarks (Bai et al., 2024; Gehrmann et al., 2023; Li et al., 2019; Zheng et al., 2024). In contrast, direct assessment can efficiently evaluate multiple items at once, but it may struggle to detect fine distinctions between outputs.
Building on (Shi et al., 2024), LLM-as-a-judge systems can be broadly categorized into two main strategies: score-based and relation-based. Similarly, (Zheng et al., 2024) proposes three variations of LLM-as-a-judge implementations that can be used independently or in combination. The first is pairwise comparison, where an LLM judge is presented with a question and two answers and tasked to determine which one is better or declare a tie. This approach can be effective for nuanced evaluations (Liu et al., 2024), but lacks scalability due to the quadratic growth of possible pairs as the number of items increases. The second is single answer grading, where an LLM judge assigns a score directly to a single answer. While this method is more scalable, it may fail to discern subtle differences between the text being evaluated, and its results may fluctuate due to changes in the judge model. The third approach is reference-guided grading, which incorporates a reference solution when applicable, enabling the model to compare answers with a gold standard. However, references are not always available, limiting its applicability. While we focus primarily on relation-based (pairwise) and score-based (direct assessment) approaches, there are several opportunities to explore nuances within these methods. For instance, our work centers on the pairwise (two-option) mode, but future research could investigate multi-option modes, such as three-option or four-option setups, which, while not altering the front-end user experience, would require adjustments in back-end calculations (Shi et al., 2024). Additionally, future exploration could include reference-guided grading, particularly in designing effective interfaces and workflows for scenarios where users have high-quality reference solutions to guide the evaluation process (Zheng et al., 2024).
Direct assessment often utilizes rubrics with multiple dimensions, while pairwise comparison may focus on a single dimension of preference. The nature of the evaluation criteria, whether objective or subjective, influences which strategy is more appropriate. For instance, subjective criteria often lack clear-cut thresholds, making comparative judgments (via pairwise) more intuitive, whereas objective criteria lend themselves to structured, rubric-based assessments. Prior work has examined how LLMs handle user-defined preferences across both objective and subjective dimensions, such as brevity, honesty, and political tone, revealing greater variability in subjective evaluations (Chiang and yi Lee, 2023; Chiang and Lee, 2023; Zheng et al., 2024). EvalAssist is the first tool that enables users to select the evaluation strategy that best fits the task, offering the flexibility to choose between direct assessment and pairwise comparison judgment strategies based on task demands and criteria type.
@@
3. Design Goals
In designing EvalAssist, we aim to address the challenges that engineers, data scientists, and researchers face when evaluating outputs and aligning criteria with their requirements. We identify five design goals to guide the development of EvalAssist, addressing key gaps in existing tools. We build on user findings from prior work with ML practitioners, which highlights the challenges faced during model evaluation (Pan et al., 2024; Desmond et al., 2024). Our approach is designed to support engineers, ML practitioners, and researchers by incorporating key features such as the ability to choose between pairwise comparison and direct assessment, (positional) bias indicators, and scalable, cost-efficient workflows. Below, we outline these goals and how EvalAssist stands apart in the evaluation tool landscape:
DG1. Isolate Generation from Evaluation. Isolating generation from evaluation is a significant need for engineers, data scientists, and practitioners working in complex multi-agent environments or retrieval-augmented generation (RAG) contexts (Gao et al., 2023), where sophisticated workflows often exist outside the evaluation tool and result in large datasets across various models, prompts, and configurations (Pan et al., 2024; Desmond et al., 2024). Other tools do not isolate the generation process from the evaluation process (Shankar et al., 2024; Arawjo et al., 2024; Kim et al., 2024a) as they focus on prompting and variations that result in output as a result of varied prompts. The targeted users for EvalAssist are developers, including engineers, ML practitioners, and researchers, who often have access to a wider variety of models and need to design complex LLM workflows to select the best model for a specific task.
DG2. Reduce Cost of Evaluation. Applying AI-assisted evaluation to large datasets typically used by research engineers, data scientists (Zheng et al., 2024) is costly and time consuming. Prior work has shown that users would like to run an evaluation on a subset of the data first (Pan et al., 2024). EvalAssist addresses the challenge of costly and time-consuming model calls by allowing users to evaluate a subset of the data, refine their criteria, and then, once satisfied, apply these criteria to the entire dataset through an SDK, a crucial feature for users handling very large datasets.
DG3. Support Multiple AI Evaluators. EvalAssist provides multiple AI evaluators. Users can select their preferred model as the evaluator, addressing issues like self-enhancement bias (Xu et al., 2024), where models favor their own responses. EvalAssist is model-agnostic, enabling users to choose any model for evaluation, unlike other front-end tools that limit users to a single AI evaluator.
DG4. Include Bias Indicators. ML researchers using LLM-as-a-judge are aware of potential biases exhibited by AI evaluation and have expressed the desire to have transparent indicators on bias in LLM-as-a-judge tooling (Pan et al., 2024). EvalAssist is the only front-end LLM-as-a-judge tool that provides a positional bias indicator. While this tool specifically focuses on positional bias, it sets the stage for addressing other issues like self-enhancement (Xu et al., 2024) and verbosity bias (Saito et al., 2023; Zheng et al., 2024) in the future. Including bias indicators helps users recognize positional biases in LLM judgments. This version explicitly flags positional bias, marking the first instance of an evaluation tool highlighting bias for users. It also paves the way for future exploration of additional biases and their impact on user responses.
DG5. Enable Flexible Evaluation Methods EvalAssist is the first tool to allow users to define their criteria for the two most common LLM-as-a-judge approaches: direct assessment (score-based, single-grading) and pairwise comparison (relation-based) (Zheng et al., 2024; Shi et al., 2024). With this tool, users can select the strategy that best suits their dataset or explore both approaches to determine which is more effective for their needs.
4. EvalAssist: System Design